A flat coil with 78 loops of wire has a radius of 15 mm. The coil is placed in a magnetic field of 0.35 T so that a maximum flux goes through it. It is later rotated in 0.07 s to a position so that no flux can go through. What is the average induced EMF produced between the terminals of the coil?


1
Expert's answer
2022-02-15T16:18:55-0500

Magnetic field

"A=\pi r^2"

"A=\pi \times0.015^2"

"e=N\frac{d\phi}{dt}"

"\phi=BAcos\theta"

"\phi_{max}=BA"

Cos0°=1


"e=78\times\frac{0.35\times\pi \times0.015^2}{0.07-0}=0.275V"